a train of length 300m crosses a cyclist going in a parallel road with a speed of 10m/s in opposite direction. how much time will it take for the train to cross the cycle if the speed of the train is 20 m/s

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

You can see the whole scenario from the frame of the train. The train sees the cyclist moving towards it with speed10m/s+20m/s=30m/s.

The distance that the cyclist must travel to cross the train=length of the train=300m

Hence the time taken for crossing=\dfrac{300m}{30m/s}=10s